# Thermodynamics of Kerr black hole: Tsallis-Cirto composition law and entropy quantization

### v6

@@ I Introduction
Hawking radiation from and Bekenstein entropy of the black holes with multiple horizons and the corresponding thermodynamics can be different from that which is obtained by consideration of only the effects of the outer horizon, see e.g. the recent paperSingha2025 and references therein. For the spherically symmetric Reissner-Nordström (RN) black hole with two horizons, several different approaches have been used.Volovik2021 ; Volovik2025TC This includes: (i) coherent Hawking radiations from two horizonz; (ii) macroscopic quantum tunneling; (iii) adiabatic transformations; (iv) singular coordinate transformations; and also (v) the Tsallis-Cirto non-extensive statistics.TsallisCirto2013 ; Tsallis2020 All these approaches demonstrate that the total entropy of the black hole is not determined by the area of the outer horizon. Due to the coherent correlations between the outer and inner horizons, the total entropy of the RN black hole and the temperature of Hawking radiation depend only on mass of the black hole and do not depend on the black hole charge.
@@ III.2 Entropy quantization for Kerr black hole
It is four times the traditionally discussed value $S_{0}=2\pi J$.
@@ III.4 Entropy quantization in Planckon toy model
Planckon model is the toy model, in which the black hole is considered as an ensemble of Planck-scale quanta, Planckons.Volovik2025P The Tsallis-Cirto $\delta=2$ statistics suggests that the entropy of the ensemble is determined by the correlations between Planckons. It is proportional to the number of the Planckon pairs, i.e. the number of ways to select two Planckons from an ensemble of $N$ Planckons:
@@
In this toy model the entropy depends on two quantum numbers, $N$ and $J$, with $J=N^{2}$ for the extremal RN black hole in the semiclassical limit $N\gg 1$. Unlike quantization by angular momentum $J$, quantization by Planckon number $N$ is the property of a toy model. The factor $4\pi$ takes place only for the Planckon mass quantum $M_{\rm P}=1/\sqrt{G}$. Nevertheless, for this mass quantum, the extreme regime is expressed entirely in integers, $J=N^{2}$, which is quite intriguing. Note that this toy model produces the results, which are similar to that of the $SU(N)$ matrix model discussed in Refs.Chu2025 ; Chu2025b , where the horizon of Schwarzschild black hole is identified with the fuzzy sphere and $N$ is the dimension of the Hilbert space of quantum gravity.
@@ III.5 Thermodynamics of rotating black holes
IV Conclusion

### v7

@@ I Introduction
Hawking radiation from and Bekenstein entropy of the black holes with multiple horizons and the corresponding thermodynamics can be different from that which is obtained by consideration of only the effects of the outer horizon, see e.g. the recent paperSingha2025 and references therein. For the spherically symmetric Reissner-Nordström (RN) black hole with two horizons, several different approaches have been used.Volovik2021 ; Volovik2025TC This includes: (i) coherent Hawking radiations from two horizons; (ii) macroscopic quantum tunneling; (iii) adiabatic transformations; (iv) singular coordinate transformations; and also (v) the Tsallis-Cirto non-extensive statistics.TsallisCirto2013 ; Tsallis2020 All these approaches demonstrate that the total entropy of the black hole is not determined by the area of the outer horizon. Due to the coherent correlations between the outer and inner horizons, the total entropy of the RN black hole and the temperature of Hawking radiation depend only on mass of the black hole and do not depend on the black hole charge.
@@ III.2 Entropy quantization for Kerr black hole
It is four times the traditionally discussed value $S_{0}=2\pi J$. Also, if Eq.(16) is written in the form
$S_{\rm Kerr\,extreme}=8\pi\sqrt{J(J+1)}\,$, (17)
this corresponds to Eq.(11) in Ref. Khriplovich1998 with an additional factor of 4. Instead of being two times smaller than the entropy of a Schwarzschild black hole of the same mass $M$, it is two times larger.
@@ III.4 Entropy quantization in Planckon toy model
Planckon model is the toy model, in which the black hole is considered as an ensemble of Planck-scale quanta, Planckons.Volovik2025P The Tsallis-Cirto $\delta=2$ statistics with the non-additive composition law in Eq.(3) suggests that the entropy of the ensemble is determined by the correlations between Planckons. It is proportional to the number of the Planckon pairs, i.e. the number of ways to select two Planckons from an ensemble of $N$ Planckons:
@@
In this toy model the entropy depends on two quantum numbers, $N$ and $J$, with $J=N^{2}$ for the extremal RN black hole in the semiclassical limit $N\gg 1$. Unlike quantization by angular momentum $J$, quantization by Planckon number $N$ is the property of a toy model. The factor $4\pi$ takes place only for the Planckon mass quantum $M_{\rm P}=1/\sqrt{G}$. Nevertheless, for this mass quantum, the extreme regime is expressed entirely in integers, $J=N^{2}$, which is quite intriguing.
Note that this toy model produces the results, which are similar to that of the $SU(N)$ matrix model discussed in Refs.Chu2025 ; Chu2025b , where the horizon of Schwarzschild black hole is identified with the fuzzy sphere and $N$ is the dimension of the Hilbert space of quantum gravity. Also, the $N^{2}$ entropy in terms of the $N\times N$ Hermitian matrices was considered for the de Sitter cosmological horizon – the dS-Matrix Theory.Susskind2023
@@ III.5 Thermodynamics of rotating black holes
IV Composition law for Schwarzschild-de Sitter black hole
It appears that the non-additive Tsallis-Cirto $\delta=2$ statistics with the composition law in Eq.(3) plays the important role in the determination of the entropy of the compact object – black hole – in Minkowski environment. If so, can this be applied to a black hole in the de Sitter environment with its cosmological horizon? In this regard, the Schwarzschild-de Sitter (SdS) black hole was considered by Susskind in his dS Matrix Theory. A discrepancy of a factor of 2 between the composition law in equation (3) and the calculated entropy deficit was found.Susskind2023 Let’s try to correct this approach.
Here we consider the composition law for the SdS entropy using approach based on the singular coordinate transformations. This approach was used in particular for the calculations of the probability of the macroscopic quantum tunneling from black hole to white hole.Volovik2022 Note that if the coordinate transformation has singularity, the diffeomorphism invariance is not necessary applicable, and the coordinate transformation may connect different physical objects. A particular example is the coordinate transformation from the black hole with mass $M$ to the white hole with the same mass, $dt\rightarrow dt+2dr\frac{v}{1-v^{2}}$, where $v^{2}=2GM/r$. This transformation has singularity at the horizon, i.e. at $v^{2}(r)=1$.
Now let us consider the SdS black hole with $v^{2}(r)=H^{2}r^{2}+\frac{2MG}{r}$, which has two horizons: the black hole horizon at $r=r_{-}$ and the cosmological horizon at $r=r_{+}>r_{-}$. Let us make the singular coordinate transformation $r\rightarrow-r$. Then one obtains the state with $v^{2}(r)=H^{2}r^{2}-\frac{2MG}{r}$, which corresponds to the negative mass of black hole. This state has only a single horizon at $r_{0}=r_{+}+r_{-}$. The entropies of the initial and final states obey the composition law:
$\sqrt{S(r_{0})}=\sqrt{S(r_{-})}+\sqrt{S(r_{+})}\,.$ (26)
This corresponds to the Tsallis-Cirto $\delta=2$ statistics and thus justifies the dS Matrix Theory suggested by Susskind.Susskind2023
Whether negative-mass objects exist in the sky is an open question.Odintsov2026 But this does not affect the consideration of the entropy composition law.
V Conclusion
@@
To further substantiate the applicability of the Tsallis-Cirto $\delta=2$ statistics to the composition law for event horizons, we considered a Schwarzschild-de Sitter black hole and demonstrated that the dS Matrix Theory proposed by SusskindSusskind2023 indeed works.